OpenStudy (anonymous):

There are (3^6)^2 ⋅ 3^0 candies in a store. What is the total number of candies in the store? 3^3 3^4 3^8 3^12

OpenStudy (anonymous):

if anyone did not know the numbers after the ^ are exponents

OpenStudy (anonymous):

is anyonehere?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[b^0=1\] always

OpenStudy (anonymous):

ok.....

OpenStudy (misty1212):

\[(3^6)^2\times 1=3^{6\times 2}\]

OpenStudy (anonymous):

any thing rest to zero is equal to 1 so here we get (3^6)^2 * 1 ====> and also if we have (x^m)^n = (x)^m*n ====> So we get (3)^6*2 = ?

OpenStudy (anonymous):

im not really understanding

OpenStudy (anonymous):

so 3^12?

OpenStudy (anonymous):

Correct !! :)

OpenStudy (anonymous):

could you help me with more?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

thank you your a life saver

OpenStudy (anonymous):

Which shows the expressions in the order they would appear on a number line from least to greatest? 17 over 9, square root of 6, square root of 15, square root of 30, 3 to the power of 3 3 to the power of 3, square root of 30, square root of 15, square root of 6, 17 over 9 3 to the power of 3, square root of 6, square root of 15, 17 over 9, square root of 30 17 over 9, square root of 30, square root of 15, 3 to the power of 3, square root of 6

OpenStudy (anonymous):

r u still here?

OpenStudy (anonymous):

Okay what is 3^3 = ?

OpenStudy (anonymous):

i really dont know im not good at algebra at all

OpenStudy (anonymous):

algebra is my worst subject

OpenStudy (anonymous):

Means cube of 3 . for example 3^2 means = 3 * 3 = 9 So 3^3 = 3*3*3 = ?

OpenStudy (anonymous):

27?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

ok....

OpenStudy (anonymous):

Similarly If we divide 17/9 =? just divide 17 by 9 what do u get ?

OpenStudy (anonymous):

1.88888889

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

ok.... then what?

OpenStudy (anonymous):

Similarly find values of square root of 6, square root of 15, square root of 30

OpenStudy (anonymous):

that i have no idea how to do

OpenStudy (anonymous):

i think the answer is a

OpenStudy (anonymous):

im not 100 percent though

OpenStudy (anonymous):

Okay here we just find least value to greatest so we know that Since 6 less than 15 is less than 30 that is 6<15<30 This implies quare root of 6< square root of 15 < square root of 30 right ?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

we know that (square root of 4 = 2) right ? So (square root of 6 = greater than 2 = 2.***) okay ?

OpenStudy (anonymous):

it means Square root of 6 is greater than 17/9 right ?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

So we have got 1.88 < quare root of 6< square root of 15 < square root of 30 < 27 So if we write expressions in the order they would appear on a number line from least to greatest we will get 17 over 9 < square root of 6 < square root of 15 < square root of 30< 3 to the power of 3.

OpenStudy (anonymous):

ok can you help me with mroe they shouldnt be like this one

OpenStudy (anonymous):

Did u understand above problem?

OpenStudy (anonymous):

ya i i said it was a a little while ago

OpenStudy (anonymous):

The amounts of food waste generated in a region during two years were 37,500,000 tons and 30,400,000 tons. What was the total food waste generated over these two years in the region, expressed in scientific notation? 67.9 ⋅ 10^14 tons 6.79 ⋅ 10^7 tons 67.9 ⋅ 10^7 tons 6.79 ⋅ 10^14 tons